OpenStudy (anonymous):

The length of a rectangle is 1 yd more than twice its width, and the area of the rectangle is 45 yd^2 . Find the dimensions of the rectangle.

OpenStudy (hoblos):

l=2w+1 A = l*w = (2w+1)(w) 45= 2w^2 +w 2w^2 +w -45 =0 (2w-9)(w+5)=0 2w-9=0 or w+5=0 w=4.5 w=-5 but w is always positive then w=4.5 yd l=2w+1 = 10 yd

OpenStudy (anonymous):

ur a genius! :-) thanks!!

OpenStudy (hoblos):

my pleasure =)